- a charge +Q at a distance 2d from a charge -Q and a point X at a distance d from -Q what is the field strength at X numerically using SI units.


1
Expert's answer
2021-10-10T16:05:23-0400
"E=k\frac{Q}{d^2}-k\frac{Q}{d^2}=0\frac{V}{m}"